# Reference Points, Risk-Taking Behavior, and Competitive Outcomes in Sequential Settings

### v3

@@ 5.3 Heterogeneity
Tables 7 and 8 show how responses to pressure vary by gender, experience, and historical rivalry frequency. Male lifters respond more strongly to both lower- and higher-ranked rivals (0.104 and 0.455, respectively) than female lifters (0.068 and 0.257), particularly under upward pressure. More experienced lifters are more responsive to higher-ranked rivals, highlighting the role of accumulated competitive cues. A higher historical rivalry frequency—defined as the number of prior encounters with a rival—is associated with greater responsiveness, suggesting that familiarity influences attempt selection. These trends indicate that male, experienced lifters facing well-known rivals take greater risks.
@@
(Pressure, lower rival)$\times$ (Historical rivalry freq) 0.100 0.127
@@
(Pressure, higher rival)$\times$ (Historical rivalry freq) 0.095 0.095
@@
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking as in the main specification. The standard errors in the brackets are clustered at federation categories.
@@
1(Turned around)$\times$ (Historical rivalry freq) 0.012 0.028
@@
1(Turning around)$\times$ (Historical rivalry freq) -0.002 -0.007
@@ IV X X X X X X
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking as in the main specification. The standard errors in the brackets are clustered at federation categories.
@@ 6 Counterfactual
Using the estimated results, we first simulate the attempt weights and the success probability for each attempt of each player under actual pressure. [^14] Our regression results use the predicted attempt weight minus the current best as a dependent variable. We obtain simulated attempt weight by adding the current best to the predicted outcomes. We then calculate the expected achieved weights, defined as the product of the predicted attempt weight and the success probability. These serve as benchmark outcomes under actual pressure conditions.
The first alternative competition setting assumes that players would have to simultaneously submit their planned attempt weights for all attempts before the first attempt and commit to the plan. This setting eliminates the possibility of pressure from rivals influencing weight attempts. While introducing this as a formal rule is unrealistic, such a commitment is common at the individual level. We refer to this scenario as the “no pressure during attempt” setting.
The second alternative competition setting posits that players could not observe the attempt weights and outcomes of their rivals after the competition begins. This removes the influence of sequentially updated outcomes from rivals on the success probability. Similar to the first scenario, formalizing this rule would be challenging, though such a commitment is feasible at the individual level. We refer to this scenario as the “no pressure during lifting” setting.
@@
For these alternative competition settings, we simulate the attempt weights, the success probability, and the expected achieved weight for each player, holding the estimated coefficients of the pressure variables at zero.
@@ 6.2 Expected weight
Figure 5 presents the distribution of expected achieved weight differences for second and third attempts, comparing the conditions of “no pressure” during the attempt, lifting, or both.
In the second attempt, almost all counterfactual settings lead to lower expected outcomes, especially in the “no pressure in both” condition, where the distribution is clearly left-skewed. The “no pressure during lifting” case shows a sharper peak near zero, suggesting that many lifters would still achieve similar weights without pressure during execution.
In the third attempt, the distributions become more dispersed, particularly under “no pressure during attempt” and “no pressure in both” conditions. A small group under “no pressure during lifting” shows gains, but most lifters experience reduced outcomes. These patterns highlight substantial heterogeneity in pressure responses and reinforce the motivational role of real-time competition.
As practical advice, most lifters may benefit from adjusting their attempts in response to competitive pressure, as removing pressure tends to reduce both attempted and achieved expected weights. However, reacting to pressure from rivals is not universally optimal across lifters. More broadly, individuals in sequential high-stakes settings should account for their own sensitivity to competitive pressure, recognizing that for many, performance deteriorates in its absence.
@@ 7 Robustness check
7.1 Choice of lifter’s current best score
As our main outcome includes the lifter’s current best, we need to take care of the construction. For example, lifters might consider the last 12 months best score as the current best and ignore the more past best score because older personal bests may not accurately reflect lifters’ current abilities. Although we partly take care about the issues by controlling age-class, we need to check robustness of our reference point choice. We consider personal best over the more recent competition, that is, the personal best achieved within a certain recent period (e.g., the last 12 months).
Table 9: Regression of the Difference between the Second and Third Attempt Weight Compared to the Current Best within a Year on Pressure
@@ 7.1 Choice of lifter’s current best score
Pressure, lower rival, 2nd 0.092
Pressure, higher rival, 2nd 0.483
Pressure, lower rival, 3rd 0.043
Pressure, higher rival, 3rd 0.427
@@
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. The standard errors in the brackets are clustered at federation categories. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from lower rival at 2nd is (lower rival 2nd attempt) - (his 1st outcome). Pressure from lower rival at 3rd is (lower rival 3rd attempt) - (his 2nd outcome). Pressure from higher rival at 2nd is (higher rival 2nd attempt) - (his 1st outcome). Pressure from higher rival at 3rd is (higher rival 3rd attempt) - (his 2nd outcome).
Table 10: Regression of Success Probability of Second and Third Attempts on Pressure, the Current Best within a Year
@@ IV X X
Notes: We control gender, body weight, equipment category, age-class, division, weight-class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. The standard errors in the brackets are clustered at federation categories.
Table 9 corresponds to Table 5 for the analysis of attempt weights using the lifter’s current best within one year. The main findings remain robust: pressure from both lower- and higher-ranked rivals continues to have positive and significant effects on attempt weights in both the second and third attempts. While the coefficient on the male dummy changes sign in the second attempt, the overall patterns are consistent, reinforcing that lifters respond to competitive pressure, particularly from higher-ranked rivals.
Similarly, Table 10 corresponds to Table 6 for the analysis of success probabilities, utilizing the lifter’s current best within a year. The main results are largely consistent with the primary specification. Upward pressure—especially attempts to overtake higher-ranked rivals—is associated with a higher likelihood of failure.
Overall, the alternative specification using the current best within a year confirms the key findings from the main analysis: competitive pressure affects both strategic weight selection and execution success, and these effects persist across time horizons.
@@ 7.2 Expected rivals’ attempt weights
In our theoretical prediction in Section 3 and empirical exercises, we use the predicted attempt weight of lifter $i$’s higher-ranked rival—unobservable to lifter $i$ at the attempt stage. Although this approach involves prediction errors, assuming that the higher rival’s attempt weight is completely unknown would ignore the pressure exerted by the higher-ranked rival, which plays a significant role in actual competition.
@@
Notes: All models include gender, body weight, equipment category, age class, division, weight class, federation, a dummy for first participation, number of competition experiences, and current best as predictors. They also include a dummy variable indicating whether the lifter shares the same declared attempt weight with another competitor—potentially causing ambiguity in the lifting order due to unresolved tie-breaking rules—and another dummy variable capturing mismatches between the actual lifting order and the interim ranking order, which may affect the set of observable information about rivals. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of the 2nd attempts are used, while in the “without 2nd success” model, the success of the 2nd attempt is excluded. Predictive accuracy is assessed using 5-fold cross-validation.
Table 11 presents the Root Mean Squared Error (RMSE) and $R^{2}$ values for different prediction models of attempt weights using linear regression. Predictive accuracy is assessed through 5-fold cross-validation. All models include gender, body weight, equipment category, age class, division, weight class, federation, a dummy for first participation, number of competition experiences, and current best as predictors. They also include a dummy variable indicating whether the lifter shares the same declared attempt weight with another competitor—potentially causing ambiguity in the lifting order due to unresolved tie-breaking rules—and another dummy variable capturing mismatches between the actual lifting order and the interim ranking order, which may affect the set of observable information about rivals. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt—unobservable when predicting the higher rival’s 2nd attempt weight—is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of both the 1st and 2nd attempts are used, whereas in the “without 2nd success” model, the success of the 2nd attempt—unobservable when predicting the higher rival’s 3rd attempt weight—is excluded.
@@
Table 12: Regression of the Difference between the Second and Third Attempt Weight Compared to the Current Best on Pressure with Bootstrap
@@
Pressure, lower rival, 2nd 0.111
Pressure, higher rival, 2nd 0.440
Pressure, lower rival, 3rd 0.068
Pressure, higher rival, 3rd 0.407
@@
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from a lower rival at the 2nd attempt is calculated as (lower rival 2nd attempt) - (his 1st outcome), and pressure from a lower rival at the 3rd attempt is (lower rival 3rd attempt) - (his 2nd outcome). For higher rivals, predicted attempt weights are obtained as the mean from the bootstrap procedure, while standard errors are computed as the standard deviation of the bootstrap estimates, ensuring that the uncertainty in the prediction process is fully reflected in the inference.
@@
To address the prediction error problem of using predicted values as explanatory variables, we implement a bootstrap procedure when incorporating the predicted attempt weight of lifter $i$’s higher-ranked rival—unobservable to lifter $i$ at the $k$-th attempt stage—into our regression framework. Since this predicted value, denoted by $\hat{Z}_{it}^{k}$, is derived from a separate estimation and contains prediction error, standard errors tend to be downward biased due to the smoothing effect of prediction. To correct for this, we apply a bootstrap approach that resamples the original data at the federation level, re-estimates the prediction model for $\hat{Z}_{it}^{k}$ in each bootstrap iteration, and then regresses the lifter’s own attempt weight or success probability on the re-estimated $\hat{Z}_{it}^{k}$. This procedure ensures that the uncertainty in $\hat{Z}_{it}^{k}$ is incorporated into the standard error estimates by reflecting its sampling variability. Consequently, the bootstrap-corrected standard errors provide more robust inference on the effect of competitive pressure from higher-ranked rivals on weight selection and performance outcomes.
@@
Compared to Table 5, which uses predicted higher rivals’ attempt weights, Table 12 incorporates a bootstrap procedure to account for uncertainty in predicted higher rival attempt weights. This leads to slightly larger standard errors but consistent point estimates and signs. Notably, the male coefficient in the second attempt shifts from –1.279 to –1.477, and pressure from higher-ranked rivals in the second attempt is stable, reinforcing the robustness of competitive pressure effects. Similarly, the effect of pressure from higher-ranked rivals in the third attempt remains stable (0.407 vs. 0.409 in Table 5). The stability of estimates confirms the validity of the benchmark results even when accounting for prediction uncertainty.

### v4

@@ 5.3 Heterogeneity
Tables 6 and 7 show how responses to pressure vary by gender, experience, and historical rivalry frequency. Male lifters respond more strongly to both lower- and higher-ranked rivals (0.104 and 0.455, respectively) than female lifters (0.068 and 0.257), particularly under upward pressure. More experienced lifters are more responsive to higher-ranked rivals, highlighting the role of accumulated competitive cues. A higher historical rivalry frequency—defined as the number of prior encounters with a rival—is associated with greater responsiveness, suggesting that familiarity influences attempt selection. These trends indicate that male, experienced lifters facing well-known rivals take greater risks.
@@
[…]
@@
[…]
@@
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking as in the main specification. Standard errors in brackets are clustered at the federation level. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
@@
[…]
@@
[…]
@@ IV X X X X X X
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking as in the main specification. Standard errors in brackets are clustered at the federation level. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
@@ 6 Counterfactual
Using the estimated results, we first simulate attempt weights and success probabilities for each attempt of each lifter under actual pressure. [^16] Our regression results use the predicted attempt weight minus the personal best as a dependent variable. We obtain simulated attempt weight by adding the personal best to the predicted outcomes. We then calculate expected achieved weights, defined as the product of the predicted attempt weight and the success probability. These serve as benchmark outcomes under actual pressure conditions.
The first alternative competition setting assumes that lifters would have to simultaneously submit their planned attempt weights for all attempts before the first attempt and commit to the plan. This setting eliminates the possibility of pressure from rivals influencing weight attempts. While introducing this as a formal rule is unrealistic, such a commitment is common at the individual level. We refer to this scenario as the “no pressure during attempt” setting.
The second alternative competition setting posits that lifters could not observe the attempt weights and outcomes of their rivals after the competition begins. This removes the influence of sequentially updated outcomes from rivals on success probability. Similar to the first scenario, formalizing this rule would be challenging, though such a commitment is feasible at the individual level. We refer to this scenario as the “no pressure during lifting” setting.
@@
For these alternative competition settings, we simulate attempt weights, success probabilities, and expected achieved weights for each lifter, holding the estimated coefficients on the pressure variables at zero.
@@ 6.2 Expected weight
Figure 5 presents the distribution of expected achieved weight differences for second and third attempts, comparing the conditions of “no pressure” during the attempt stage, during the lifting stage, or during both stages.
In the second attempt, most counterfactual outcomes are negative, especially in the “no pressure during both stages” condition, where the distribution is clearly left-skewed. The “no pressure during lifting” case shows a sharper peak near zero, with only limited gains on the positive side.
In the third attempt, the negative side remains dominant, particularly under “no pressure during attempt” and “no pressure during both stages” conditions. The “no pressure during lifting” case still includes some positive outcomes, but most observations are at or below zero. These patterns highlight substantial heterogeneity in pressure responses and reinforce the motivational role of real-time competition.
As practical advice, most lifters may benefit from adjusting their attempts in response to competitive pressure, as removing pressure tends to reduce both attempt weights and expected achieved weights. However, reacting to pressure from rivals is not universally optimal across lifters. More broadly, individuals in sequential high-stakes settings should account for their own sensitivity to competitive pressure, recognizing that for many, performance deteriorates in its absence.
@@ 7 Robustness check
7.1 Alternative personal best definition
Our main specification uses each lifter’s all-time personal best as an ability-based baseline for measuring deviations in attempt weight. However, older records may not accurately reflect a lifter’s current ability, so we check robustness by replacing the all-time personal best used in the dependent variable with the trailing twelve-month personal best—the highest outcome achieved within the 12 months preceding the competition. The pressure variables remain unchanged, as they are defined in terms of rivals’ declared or predicted attempt weights and the focal lifter’s realized best outcome within the competition.
Table 8: Regression of the Difference between Attempt Weight and Trailing Twelve-Month Personal Best on Pressure
@@ 7.1 Alternative personal best definition
Pressure, lower rival, 2nd 0.092***
Pressure, higher rival, 2nd 0.483***
Pressure, lower rival, 3rd 0.043***
Pressure, higher rival, 3rd 0.427***
@@
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. Standard errors in brackets are clustered at the federation level. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from the lower rival at the second attempt is (lower rival second-attempt weight) - (focal lifter’s current best outcome after the first attempt). Pressure from the lower rival at the third attempt is (lower rival third-attempt weight) - (focal lifter’s current best outcome after the second attempt). Pressure from the higher rival at the second attempt is (predicted higher rival second-attempt weight) - (focal lifter’s current best outcome after the first attempt). Pressure from the higher rival at the third attempt is (predicted higher rival third-attempt weight) - (focal lifter’s current best outcome after the second attempt). $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
Table 9: Regression of Success Probability of Second and Third Attempts on Pressure, Trailing Twelve-Month Personal Best
@@ IV X X
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. Standard errors in brackets are clustered at the federation level. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
Table 8 corresponds to Table 4 for the analysis of attempt weights using the trailing twelve-month personal best. The main findings remain robust: pressure from both lower- and higher-ranked rivals continues to have positive and significant effects on attempt weights in both the second and third attempts. While the coefficient on the male dummy changes sign in the second attempt, the overall patterns are consistent, reinforcing that lifters respond to competitive pressure, particularly from higher-ranked rivals.
Similarly, Table 9 corresponds to Table 5 for the analysis of success probabilities, utilizing the trailing twelve-month personal best. The main results are largely consistent with the primary specification. Upward pressure—especially attempts to overtake higher-ranked rivals—is associated with a higher likelihood of failure.
Overall, the alternative specification using the trailing twelve-month personal best confirms the key findings from the main analysis: competitive pressure affects both strategic weight selection and execution success, and these effects persist across time horizons.
@@ 7.2 Expected rivals’ attempt weights
In our empirical exercises, we use the predicted attempt weight of lifter $i$’s higher-ranked rival—unobservable to lifter $i$ at the attempt stage. Although this approach involves prediction errors, assuming that the higher rival’s attempt weight is completely unknown would ignore the pressure exerted by the higher-ranked rival, which plays a significant role in actual competition.
@@
Notes: All models include gender, body weight, equipment category, age class, division, weight class, federation, a dummy for first participation, number of competition experiences, and personal best as predictors. They also include a dummy variable indicating whether the lifter shares the same declared attempt weight with another competitor—potentially causing ambiguity in the lifting order due to unresolved tie-breaking rules—and another dummy variable capturing mismatches between the actual lifting order and the interim ranking order, which may affect the set of observable information about rivals. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of the 2nd attempt are used, while in the “without 2nd success” model, the success of the 2nd attempt is excluded. Predictive accuracy is assessed using 5-fold cross-validation.
Table 10 presents the Root Mean Squared Error (RMSE) and $R^{2}$ values for different prediction models of attempt weights using linear regression. Predictive accuracy is assessed through 5-fold cross-validation. All models include gender, body weight, equipment category, age class, division, weight class, federation, a dummy for first participation, number of competition experiences, and personal best as predictors. They also include a dummy variable indicating whether the lifter shares the same declared attempt weight with another competitor—potentially causing ambiguity in the lifting order due to unresolved tie-breaking rules—and another dummy variable capturing mismatches between the actual lifting order and the interim ranking order, which may affect the set of observable information about rivals. In the full model for the 2nd attempt, both the attempt weight and the success of the 1st attempt are included as predictors, while in the “without 1st success” model, the success of the 1st attempt—unobservable when predicting the higher rival’s 2nd attempt weight—is excluded. Similarly, in the full model for the 3rd attempt, the attempt weight and the success of both the 1st and 2nd attempts are used, whereas in the “without 2nd success” model, the success of the 2nd attempt—unobservable when predicting the higher rival’s 3rd attempt weight—is excluded.
@@
Table 11: Regression of the Difference between the Second and Third Attempt Weight Compared to the Personal Best on Pressure with Bootstrap
@@
Pressure, lower rival, 2nd 0.111***
Pressure, higher rival, 2nd 0.440***
Pressure, lower rival, 3rd 0.068***
Pressure, higher rival, 3rd 0.407***
@@
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. $W_{it}^{2}$ and $W_{it}^{3}$ are lifter $i$’s attempt weights in the second and third attempts in competition $t$. Pressure from the lower rival at the second attempt is (lower rival second-attempt weight) - (focal lifter’s current best outcome after the first attempt). Pressure from the lower rival at the third attempt is (lower rival third-attempt weight) - (focal lifter’s current best outcome after the second attempt). Pressure from the higher rival at the second attempt is (predicted higher rival second-attempt weight) - (focal lifter’s current best outcome after the first attempt). Pressure from the higher rival at the third attempt is (predicted higher rival third-attempt weight) - (focal lifter’s current best outcome after the second attempt). For higher rivals, predicted attempt weights are obtained as the mean from the bootstrap procedure, while standard errors are computed as the standard deviation of the bootstrap estimates, ensuring that the uncertainty in the prediction process is fully reflected in the inference. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
@@
To address the prediction error problem of using predicted values as explanatory variables, we implement a bootstrap procedure when incorporating the predicted attempt weight of lifter $i$’s higher-ranked rival—unobservable to lifter $i$ at the $k$-th attempt stage—into our regression framework. Since this predicted value, denoted by $\hat{Z}_{it}^{k}$, is derived from a separate estimation and contains prediction error, standard errors tend to be downward biased due to the smoothing effect of prediction. To correct for this, we apply a bootstrap approach that resamples the original data at the federation level, re-estimates the prediction model for $\hat{Z}_{it}^{k}$ in each bootstrap iteration, and then regresses the lifter’s own attempt weight or success probability on the re-estimated $\hat{Z}_{it}^{k}$. This procedure ensures that the uncertainty in $\hat{Z}_{it}^{k}$ is incorporated into second-stage inference by reflecting its sampling variability.
@@
Compared to Table 4, which uses predicted higher rivals’ attempt weights, Table 11 incorporates a bootstrap procedure to account for uncertainty in predicted higher rival attempt weights. Standard errors are broadly similar in magnitude (some coefficients are larger and others are smaller), while point estimates and signs remain consistent. Notably, the male coefficient in the second attempt shifts from –1.279 to –1.477, and pressure from higher-ranked rivals in the second attempt is stable, reinforcing the robustness of competitive pressure effects. Similarly, the effect of pressure from higher-ranked rivals in the third attempt remains stable (0.407 vs. 0.409 in Table 4). The stability of estimates confirms the validity of the benchmark results even when accounting for prediction uncertainty.
7.3 Alternative definitions of risk-taking
We also examine whether our results are robust to an alternative scaling of the dependent variable. Instead of the level difference $W_{it}^{k}-\text{(best)}$, we rescale this difference by the within-weight-class standard deviation of personal bests, yielding $(W_{it}^{k}-\text{(best)})/\text{sd}(\text{best})$. Because the dispersion of personal bests differs across weight classes, the same absolute deviation may reflect a different degree of risk-taking in classes with substantially different ability levels. In our regression specification, weight-class fixed effects absorb systematic level differences across classes, while this rescaling adjusts for differences in the dispersion of personal bests within classes. Tables 12 and 13 report the results. Column (1) reproduces the baseline specification and Column (2) uses the standardized difference. The coefficients on pressure from both lower- and higher-ranked rivals remain positive and statistically significant across both specifications and for both attempts. The $R^{2}$ values are comparable between the two specifications, and the qualitative conclusions are unchanged: competitive pressure significantly increases attempt weight selection under both scalings.
Table 12: Robustness: Alternative Dependent Variable, Second Attempt
(1) (2)
Dependent Variable $W_{it}^{2}$-(best) $(W_{it}^{2}$-(best))/sd(best)
Male -1.254*** 0.012
(0.460) (0.012)
Body weight 0.267*** 0.003***
(0.025) (0.000)
Num experience 0.166 0.001
(0.115) (0.002)
1(first participation) 31.479*** 0.547***
(1.372) (0.025)
Pressure, lower rival 0.101*** 0.001***
(0.008) (0.000)
Pressure, higher rival 0.425*** 0.007***
(0.012) (0.000)
Control X X
Num.Obs. 204215 204217
R2 0.514 0.495
R2 Adj. 0.512 0.493
Notes: Column (1) uses the level difference $W_{it}^{2}-\text{(best)}$ as the dependent variable. Column (2) standardizes the level difference by the within-weight-class standard deviation of personal bests: $(W_{it}^{2}-\text{(best)})/\text{sd}(\text{best})$. We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
Table 13: Robustness: Alternative Dependent Variable, Third Attempt
(1) (2)
Dependent Variable $W_{it}^{3}$-(best) $(W_{it}^{3}$-(best))/sd(best)
Male 3.439*** 0.103***
(0.633) (0.014)
Body weight 0.347*** 0.005***
(0.049) (0.001)
Num experience 0.178 0.001
(0.149) (0.002)
1(first participation) 35.362*** 0.608***
(1.821) (0.032)
Pressure, lower rival 0.071*** 0.001***
(0.008) (0.000)
Pressure, higher rival 0.415*** 0.007***
(0.012) (0.000)
Control X X
Num.Obs. 181232 181233
R2 0.369 0.352
R2 Adj. 0.366 0.350
Notes: Column (1) uses the level difference $W_{it}^{3}-\text{(best)}$ as the dependent variable. Column (2) standardizes the level difference by the within-weight-class standard deviation of personal bests: $(W_{it}^{3}-\text{(best)})/\text{sd}(\text{best})$. We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.